# A pure dual approach for hedging Bermudan options

### v1

@@ 4.2.1 Max-call option
We have indicated in Figure 3 the P&L associated with the corresponding hedging strategies. Again, we note a clear interest to include European options in the hedging portfolio. The P&L variance is reduced from 7.3 to 1.5. We note that the P&L distribution of the stock only portfolio is slightly shifted to the right. This is expected since its mean value is exactly the difference between $\hat{U}_{0}^{Q}$ and the Longstaff Schwartz price, which is equal to $8.55-8.1=0.45$ on this example.
Refer to captionRefer to caption
Figure 3: P&L histograms of the hedging strategy for the Bermudan max-call option of Table 3 obtained with $\bar{N}=5$, $P=10$, $Q=2\times 10^{6}$ for the stock only strategy (left) and the strategy using extra European options (right).
@@ 4.2.3 Basket option
[…]
100000 1 False 4.42 4.43
100000 1 True 4.25 4.28
100000 5 False 4.19 4.30
100000 5 True 4.02 4.23
500000 5 False 4.26 4.28
500000 5 True 4.15 4.19
1000000 10 False 4.24 4.26
1000000 10 True 4.14 4.18
Table 9: Prices for a put option on $3$-dimensional basket using a polynomial approximation of order $5$ with $K=S_{0}=100$, $T=1$, $r=0.05$, $\sigma^{i}=0.2$, $\rho=0.3$ and $10$ exercising dates. The Longstaff Schwartz algorithm with a polynomial approximation of order $3$ gives $4.03$.
Refer to captionRefer to caption

### v2

@@ 4.2.1 Max-call option
We have indicated in Figure 3 the P&L associated with the corresponding hedging strategies. Again, we note a clear interest to include European options in the hedging portfolio. The P&L variance is reduced from 7.3 (left plot) to 1.5 (center plot). We note that the P&L distribution of the stock only portfolio is slightly shifted to the right. This is expected since its mean value is exactly the difference between $\hat{U}_{0}^{Q}$ and the Longstaff Schwartz price, which is equal to $8.55-8.1=0.45$ on this example. We have also plotted, on the right of Figure 3, the P&L obtained with the Monte Carlo method proposed by Wang and Caflisch (2010), when the delta hedging strategy is calculated once and for all at the initial time. Even if the magnitudes of the hedging errors are of the same order, the left plot is more peaked than the right one, which indicates a better hedge provided by our algorithm. Besides this, we believe that one strength of our algorithm is its ability to add any available financial instruments in the hedging portfolio. Of course, Monte Carlo methods such as the one of Wang and Caflisch (2010) give also approximations of the Gamma matrix, but it is not always obvious to translate this matrix in terms of financial instruments (European options on single assets or on many assets with different strikes). Our algorithm gets around this difficulty and directly computes the quantity of financial instruments for the hedging strategy.
Figure 3: P&L histograms of the hedging strategy for the Bermudan max-call option of Table 3 obtained with $\bar{N}=5$, $P=10$, $Q=2\times 10^{6}$ for the stock only strategy (left) and the strategy using extra European options (center). The right plot shows the result obtained with the delta hedging strategy calculated with Wang and Caflisch (2010).
@@ 4.2.3 Basket option
Data availability statement: Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Conflict of interest: The authors have no conflict of interest to declare.